OpenStudy (anonymous):

I really need help on this one question its a 8th grade one, so maybe its easy for any of you...

OpenStudy (anonymous):

y=-4x^2-8x-14 x=-3 I HAVE TO SOLVE AND FIND A WAY FOR IT TO EQUAL -26

OpenStudy (anonymous):

oops...caps

OpenStudy (anonymous):

Ohh we are so offended by your caps

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

I'm having the same exact trouble as amonoconnor XD

OpenStudy (anonymous):

If you simply replace x by -3 and solve it, you get -26

OpenStudy (anonymous):

\[y=-4*(-3)^2-8*(-3)-14=-26\]

OpenStudy (anonymous):

I did work it out...im doing something wrong and I don't know what...I keep getting -46...

OpenStudy (anonymous):

Lets take it in parts, shall we?

OpenStudy (anonymous):

\[-4*(-3)^2\] What is this?

OpenStudy (anonymous):

I believe its -36...

OpenStudy (anonymous):

Yes it is. It is -36. And next \[-8*(-3)=?\]

OpenStudy (anonymous):

didnt see square

OpenStudy (anonymous):

but that's minus eight...its not suppose to be a negative right?

OpenStudy (anonymous):

It is a negative. So \[-8*(-3)=24\]

OpenStudy (anonymous):

Because you had \[-8x\] in the equation

OpenStudy (anonymous):

but its actually 4x^-(8x)

OpenStudy (anonymous):

So this gives us: \[-4*(-3)^2=-36\] \[-8*(-3)=24\] and \[-14\] So: \[y=-36+24-14=?\]

OpenStudy (anonymous):

It is \[y=-4x^2-8x-14\] right?

OpenStudy (anonymous):

so that's what I was doing wrong XD

OpenStudy (anonymous):

thanks, I finally get it

OpenStudy (anonymous):

You're welcome :)